the area of triangle with base, b and height h, is given by:
A =1/2bh
if the area of a certain triangle is 38 square centimeters and the base is 8 centimeters, what is the height? please show all of your work.

Answer by stanbon(75887)

A =1/2bh
if the area of a certain triangle is 38 square centimeters and the base is 8 centimeters, what is the height?
------
38 = (1/2)*8*h
---
76 = 8h
---
h = 76/8
===
height = 9.5 cm
======================
Cheers,
Stan H.
======================